Write an introduction about performance characteristics of heat pump Write clearly
Answer #1

INTRODUCTION

  • A heat pump is a device that extracts thermal energy from a low-temperature source (such as the outside air or ground) and transfers it to a higher-temperature sink (such as the heated indoors of a building).
  • For instance, a household refrigerator may be thought of as a heat pump. While cooling the food products, it takes thermal energy from them. The outside coil, on the other hand, transfers thermal energy to the ambient air. The food acts as the source, the air in the kitchen as the sink.
  • Basic laws in thermodynamics recognize that work is not required to transfer heat from a high-temperature source to a lower-temperature sink (a simple conductive path would suffice), but it is required to transfer thermal energy from a low-temperature source to a higher-temperature sink.
  • The shaft work per unit time, It is the input needed to operate the heat pump, while TH is the rate of thermal energy transferred to the warmer environment, and TC , is that taken from the colder environment. The case represented is that of heating the warmer indoors, at TH while extracting thermal energy from the colder outdoors, at TC.
  • The desired output in this case of a heating mode is TH. Hence, the coefficient of performance of the heat pump, in the heating mode, would be :

COP = QH /W =TH/W =T​​H/ TH - TC

Where - TH is higher inside temperature

T​​​​​​C is lower outside temperature

W= Q​​​​​​H - QC = TH - TC

QH - higher heat rate

QL - lower heat rate

Hence, COP > 1.0.